the sum of two integers is ten, three times the larger integer is three less than eight times the smaller integer. find de integer
Answer by jorel1380(3719)

m+n=10
8m-3=3n
10-m=n
8m-3=3(10-m)
8m-3=30-3m
11m=33
m=3
n=7
The two integers are 3 and 7..